### v2

# The effects of fibre spatial distribution and relative orientation on the percolation and mechanics of stochastic fibre networks: A model of peptide hydrogels

Abstract
The structures of fibre networks can vary greatly due to fibre interactions during formation. We have modified the steps of generating Mikado networks to create two new model classes by altering the spatial distribution and relative orientation of their fibres to mimic the structures of self-assembling peptide hydrogels (SAPHs), whose physical properties depend strongly on their fibres’ interactions. The results of our models and experiments on a set of beta-sheet forming SAPHs show that modifying a network’s structure affects the percolation threshold and the mechanical behaviour of the material, both near percolation and at higher densities.
Many materials are constructed from fibre networks, from biological materials, such as collagen-based tissues and the cytoskeleton, to synthetic materials, such as gels and paper [1]. These materials are often modelled as stochastic networks [2, 3, 4, 5] to study their physical properties, such as percolation [6] and elastic modulus [2, 7]. It has been shown that, at densities near percolation, the network’s elastic moduli vary proportionally to $(\rho-\rho_{c})^{f}$, where $\rho$, $\rho_{c}$, and $f$ are the density ($\rho=N/L^{2}$, the number of fibres per unit square area of side length $L$), percolation threshold density, and elastic exponent of the network, respectively [8]. At higher densities ($\rho\gg\rho_{c}$), they vary proportionally to $\rho^{f_{h}}$, where $f_{h}$ is the high density elastic exponent [9, 10].
Researchers have modified various aspects of fibre networks to investigate their material properties. Parvez and Picu [11] investigated network connectivity – the mean number of fibre segments emerging from each node in the network (a segment is the portion of a fibre between two connected nodes). They showed that the elastic modulus in the non-affine regime increases, but the exponent $f_{h}$ decreases, with increasing connectivity. The global orientation preference of fibres has also attracted attention [12, 13]. Tarasevich and Eserkepov [12] modified fibre orientation with respect to a global coordinate system to create anisotropic networks and showed that they percolate at higher densities.
Network-based materials form in a wide variety of ways [14], which gives them dissimilar structures. For example, open cell foams differ from collagen gels. Foams are formed by introducing gas bubbles into monomer solution before solidification [15], while collagen gels are formed by cells laying down fibrils which self-assemble [16]. To model the former, it is more suitable to use Voronoi networks, and for the latter, Mikado networks [17].
In this paper, we model self-assembling peptide hydrogels (SAPHs), a family of peptides extensively studied by our group for their self-assembly, gelation, and potential biomedical uses like tissue engineering and drug delivery [18, 19, 20]. In SAPHs, fibres formed of short peptide sequences create a network through self-association [21]. The fibres are sufficiently thick and rigid enough to ignore thermal fluctuations. Variations in amino acid sequences or environmental conditions can significantly alter fibre interactions, affecting the network’s structure and mechanical properties, with elastic exponents ranging from 2 to 4.6 [10, 22]. Despite their importance, however, the effects of local fibre interactions on the structure and mechanical behaviour of the network are not fully understood. To understand how the structural changes caused by fibre interactions lead to mechanical changes, we modified the Mikado network generation process (we did not directly model fibre formation, branching and interaction). Our results apply equally to any materials that have similar structures.
Two dimensional Mikado networks are created by randomly and independently placing a given number of fibres of length $l$ (we used $l=1$) with random orientation within a square domain of side length $L$. We modified this process by changing two fibre features, their distribution in space and their relative orientation, as follows:
  • To modify spatial distribution, after randomly determining a position for a fibre’s centre, a probability, $P_{e}$, was assigned based on the number of existing fibre centres within a radius, $r_{e}$. A random number between 0 and 1 was then generated, and if it exceeded $P_{e}$, the fibre was deleted.
  • To modify orientation, the orientations of fibres within a radius, $r_{\theta}$, were used to define the distribution from which the candidate fibre’s orientation was drawn, increasing the likelihood of a specified alignment relative to existing fibres.
The defined probabilities and the critical radii can be specified in various forms. The forms we chose are given below. For the existence probability, we specified
$P_{e}=(1-P_{r})^{n_{e}},$ (1)
$r_{e}=\alpha\sqrt{\frac{1}{\rho}},$ (2)
where $P_{r}$, $n_{e}$, and $\alpha$ are, respectively: the reduction in the probability of the new fibre existing due to the presence of a single fibre centre previously placed within the circle, the number of fibre centres that already exist within that circle, and a constant which scales $r_{e}$ relative to $\sqrt{1/\rho}$, the distance between nodes in perfect square lattice. In our simulations we chose $\alpha=3/4$. The orientation is chosen, based on the orientations of the existing fibres within a circle of radius $r_{\theta}$ (we chose $r_{\theta}=l/2)$, from a Gaussian mixture distribution [23]:
$g(\theta)=\sum_{i=1}^{n_{\theta}}\frac{1}{n_{\theta}}\frac{1}{\sigma\sqrt{2\pi}}\exp{\left(-\frac{(\theta-(\theta_{fi}+\Theta))^{2}}{2\sigma^{2}}\right)},$ (3)
where $\Theta$, $n_{\theta}$, $\sigma$, and $\theta_{fi}$, are the preferred orientation relative to the existing fibres, the number of fibres within the circle, the standard deviation of the distribution (we used $\sigma=1/\sqrt{12}$), and the angle the $i^{\text{th}}$ fibre makes with the $x$-axis, respectively.
We generated and compared three network classes: the unmodified Mikado network, the distribution-modified (DM) network, where only spatial distribution is modified, and the distribution-orientation-modified (DOM) network, where both spatial distribution and orientation are modified. For the DOM networks, we considered two orientations: $\Theta=0$ (which encourages fibres to co-align) and $\Theta=\pi/2$ (which encourages the fibres to be perpendicular). Networks with the parameters given in Table 1 were studied, and examples of each type are shown in Figure 1.
Table 1: Parameters used to generate the networks, the critical percolation threshold density, $\rho_{c}$, elastic exponent (from equation (5)), $f$, and exponent at high density, $f_{h}$, for the six network classes considered (mean $\pm$ standard error).
Network type $P_{r}$ $\Theta$ $\rho_{c}l^{2}$ $f$ $f_{h}$
Mikado N/A N/A $5.6373\pm 0.0001$ $3.77\pm 0.02$ $7.9\pm 0.1$
DM 0.8 N/A $4.9919\pm 0.0002$ $3.56\pm 0.01$ $7.7\pm 0.1$
DOM 1 0 $\frac{\pi}{2}$ $4.954\pm 0.001$ $3.70\pm 0.01$ $8.4\pm 0.1$
DOM 2 $0.8$ $\frac{\pi}{2}$ $4.1907\pm 0.0001$ $3.53\pm 0.01$ $8.0\pm 0.1$
DOM 3 00 00 $9.43\pm 0.01$ $3.92\pm 0.02$ $9.6\pm 0.1$
DOM 4 $0.8$ 00 $7.690\pm 0.001$ $3.68\pm 0.02$ $9.0\pm 0.1$
[…]
Figure 1: Samples of the studied networks with $\rho=30$, $l=1$, and $L=5$.
We first studied percolation. All the subsequent analysis was carried out on percolated systems. Network percolation was investigated based on the finding that $\rho_{0.5}(L)$ (the density at which the percolation probability is 0.5 for finite $L$) converges to $\rho_{c}$ as $L\rightarrow\infty$ [6, 24, 25], since
$\rho_{0.5}(L)-\rho_{c}=-\frac{b_{0}}{a_{1}}L^{-1-1/\nu_{c}}+...\leavevmode\nobreak\ ,$ (4)
where $\rho_{c}$, $\nu_{c}$, $b_{0}$, and $a_{1}$ are the critical percolation density, the correlation-length exponent (a universal constant that is only affected by dimensionality, which equals $4/3$ for 2D systems [6, 8]), and some fitting constants, respectively. We used a method similar to that developed by Pike and Seager [26] to find $\rho_{0.5}(L)$ (see Supplemental Material for further details).
The percolation thresholds are presented in Table 1. Our results for Mikado networks are consistent with the literature [6, 27]; however, as we modify the spatial dispersion, or make the fibres more likely to be perpendicular, $\rho_{c}$ decreases, and if we make the fibres more likely to be parallel, $\rho_{c}$ increases.
The second aspect that we investigated was how fibres are dispersed within the domain of study. To do so, we created networks of various densities with $L=10$, divided them into 400 smaller boxes, counted the number of fibre centres in each box, then calculated the mean and variance of this quantity. It is well known that, in Mikado networks, the number of fibre centres in each box follows a Poisson distribution [28] (Figure 2(a)). An increase in the value of $P_{r}$ results in a more homogeneous (lower variance) network.
Figure 2: Geometric features of the networks. (a) The mean, $\overline{x}$, and variance, $s^{2}$, of the number of fibres in 400 sub-boxes, as described in the text, for Mikado Networks and DM networks with various values of $P_{r}$. (b) The connectivity, $Z$, as a function of the dimensionless density, $\rho l^{2}$. (c) The mean segment length, $\overline{l_{c}}$. (d) The mean number of segments per unit area, $\overline{N_{c}}/L^{2}$. The standard errors for this data are smaller than the size of the markers in all plots.
The connectivity of a network, $Z$, plays an important role in the network’s behaviour [11]. As depicted in Figure 2(b), for more homogeneous networks, the connectivity is slightly lower. The reason for this is fibre clumping (see Figure 1) [28]. As a result, fibres in the vicinity of these clumps have a relatively high chance of forming connections, and thus will have a higher connectivity. Moreover, for the networks whose fibres have a preference for being parallel ($\Theta=0$), the connectivity is considerably smaller. However, the difference between all network classes vanishes as the density increases.
Two other important geometrical factors are the mean length and density of fibre segments, as plotted in Figures 2(c) and 2(d). Networks with $\Theta=0$ have longer fibre segments and lower segment density but gradually converge with the other networks as $\rho l^{2}\rightarrow\infty$. It will be shown that this affects the modulus of the networks substantially. In contrast, other network classes converge at lower densities, suggesting relatively similar mechanical responses.
Our investigation into the networks’ mechanical properties is divided into the two cases of near percolation and high density [2, 7]. For low densities, the finite size scaling method [8, 7] is used, and at high densities, a representative volume element (RVE) is considered [9]. In this study, the segments of the network are considered to be beams with second moment of area, $I$ [29], with square cross-sections. The nodes are considered to be stiff cross-links (relative rotation is prohibited between beams connected at a node). The Timoshenko beam model [29] is used to model the beams and the direct stiffness method [29] is used to model infinitesimal deformation of the network.
It has been shown that the following scaling law can be used to describe the relationship between the mean Young’s modulus, $\overline{E}$, and the size of the network at $\rho=\rho_{c}$, [30]
$\overline{E}=c_{1}L^{-f/\nu_{c}},$ (5)
where $c_{1}$ is a fitting constant. This is equivalent to $-\frac{\ln{\overline{E}}}{\ln{L}}=\frac{f}{\nu_{c}}-\frac{\ln{c_{1}}}{\ln{L}}$. By calculating $\overline{E}$ for various system sizes, and using a plot with $-\frac{\ln{\overline{E}}}{\ln{L}}$ and $\frac{1}{\ln{L}}$ as the axes, the intersection of the fitted curve with the $-\frac{\ln{\overline{E}}}{\ln{L}}$ axis is $\frac{f}{\nu_{c}}$.
To calculate the elastic modulus, a displacement in the $y$-direction was applied to each node on the top boundary, while the displacement in the $x$-direction and rotation were fixed. On the bottom boundary, zero displacement and rotation was imposed, and on the lateral boundaries, zero force was imposed. The number of realisations varied with system size from 1,600 for the largest to 102,600 for the smallest.
The results are presented in Table 1. For the Mikado networks, the value of the elastic exponent was calculated as $f=3.73\pm 0.02$. This is different from a previously reported value, $f=3.2\pm 0.3$ [7, 2], due to our use of a more accurate value for the percolation threshold. Our results show that the various network types considered here have different elastic exponents. This suggests that materials whose fibres interact during formation, thus modifying the structure of the network, belong to different universality classes.
It has been shown that, at high density, the behaviour of fibre networks can be divided into two regimes: affine and non-affine [9, 2, 7]. In the non-affine regime, the bending stiffness of the fibre segments is much smaller than their stretching stiffness. In the affine regime, the network elastic moduli vary linearly with density and in the non-affine regime, they vary proportionally to $(\rho l)^{f_{h}}(l_{b}/l)^{2}$, where $l_{b}=\sqrt{I/A}$, and $A$ is the fibre cross-sectional area [9].
To determine the shear modulus, $G$, of the simulated networks, a periodic boundary condition with a far-field simple shear deformation was applied in a manner that is explained in depth by Nguyen et al. [31]. In order to find $f_{h}$ for different network types, we simulated networks with a fibre thickness of 0.0001 and densities ranging from 25 to 500; the results are presented in Figure 3(a). For the Mikado network, $f_{h}$ was found to be 7.9, as reported in the literature [9, 32]. The other values are presented in Table 1. Increasing the spatial dispersion of fibres slightly decreases the value of $f_{h}$. Furthermore, noticeable changes can be observed for networks with preferred parallel orientations. It is likely that all of the geometrical features discussed in Figure 2 impact the network mechanical moduli, but disentangling every specific relationship between structure and mechanics is not straightforward. However, as can be seen from Figures 2b-d, for DOM type 3 and 4 networks, the connectivity is lower, and the gradients of the mean segment length and the mean segment density are steeper; which, in the bending-dominated regime, causes the value of $f_{h}$ to be higher.
The value of $G$ varies significantly between the six network types. It can vary up to 50 times in the bending-dominated regime. For DOM type 3 and 4 networks, $G$ is smaller by at least one order of magnitude in the non-affine regime. This is due to their lower connectivity and segment density, and higher segment length. This clearly shows how influential network structure is on the mechanical behaviour of fibrous materials.
Figure 3: (a) Shear modulus as a function of density for high density networks ($E_{f}$ is the fibre Young’s modulus). (b) The dependence of Poisson’s ratio on density and $l_{b}$. For the main figure $l_{b}=2.9\mathrm{e}{-5}$. In the inset panel, $\rho=75$. (c) $PS_{n\%}$ represents the smallest possible proportion of a network’s segments that together store $n\%$ of the total energy in the network. $PS_{n\%}^{\rho=25}$ represents $PS_{n\%}$ with $\rho=25$. All the data is this figure is for Mikado networks in the non-affine regime. (d) Histogram of the orientations of fibres relative to the horizontal axis (which is perpendicular to the stretch direction) that carry $50\%$ of the energy in each Mikado network. The absolute value of each angle has been taken. This data is for $l_{b}=2.9\mathrm{e}{-6}$.
We also investigated the Poisson’s ratio, $\nu$, of our six network types. To do so, we applied boundary conditions that impose a pure longitudinal deformation, and calculated the total forces in the directions parallel and perpendicular to the direction of stretch; Poisson’s ratio is then calculated using $\nu=F_{\perp}/F_{\parallel}$, where $F_{\perp}$ and $F_{\parallel}$ are the total perpendicular and parallel forces at the boundary interfaces (see Supplemental Material for a derivation). We display the effects of changing the density, $\rho$, and $l_{b}$, in Figure 3(b). There is a noticeable difference in the value of $\nu$ between the different networks, which is more pronounced when the relative orientations of the fibres change.
The Poisson’s ratio in the non-affine regime is not constant and changes with increasing density. This is contrary to previous papers that reported $\nu$ to be constant [2, 32]. The effect on $\nu$ of increasing density diminishes into the affine regime (as far as we have made observations). The effect of $l_{b}$, however, is different; in the non-affine regime, increasing it has no effect, but as we transition into the affine regime, increasing it reduces the value of $\nu$ as far as we have made observations.
The reason for the effect of density on $\nu$, even when the network remains in the non-affine regime, lies in how the network absorbs energy at different densities. We have demonstrated this for Mikado networks by looking at two parameters: the percentage of segments that carry a given percentage of the total energy absorbed by the network, and the orientations of these fibres. The results are presented in Figures 3(c) and 3(d). The percentage of segments involved in absorbing a given portion of the total energy decreases with increasing density. The orientation distribution of these fibres also noticeably changes. In Figure 3(d), we can see that, at higher densities, the segments have a greater preference to be parallel to the stretch direction (in all cases the energy absorbed through stretching is negligible). The reason for the change in the energy distribution is unclear, but it can be inferred that this change is rooted in how the microstructure deforms and this, in turn, affects the Poisson’s ratio.
To compare the results of the model networks qualitatively with real SAPHs, we carried out shear rheometry tests on three $\beta$-sheet forming peptides: KFEFEFKFKF (F), KFEFEFKFKK (K), and KFEFEFKFKFK (FK) (K: lysine; F: phenylalanine; E: glutamic acid). These peptides are based on a design first developed by Zhang et al. [33], Zhang and Altman [34] (see the Supplemental Material for details). Figure 4a-c shows transmission electron microscopy (TEM) images obtained for 10$\times$ diluted hydrogels alongside the chemical structure of the SAPHs. These images confirm the formation in all three systems of semi-rigid peptide fibre networks with the thinnest fibre diameter ranging from 3 to 4 nm. Wychowaniec et al. [22] have shown that peptides with hydrophobic end-groups (F) have more hydrophobic edge interactions and are thought, therefore, to form more crosslinks. These networks are thus more similar to the perpendicular network model described above. On the other hand, peptides with hydrophilic (K) end-groups were shown to make fewer crosslinks due to edge electrostatic repulsion and, therefore, will be more similar to the parallel network model. In Figure 4d, the storage moduli, $G^{\prime}$, obtained for these three hydrogels are presented. For FK and K hydrogels, similar exponents (2.4 $\pm$ 0.2) were obtained. Both these systems have fibres with hydrophilic edges due to the presence of the terminal K. FK was found to have a higher $G^{\prime}$ compared to K which can be attributed to the presence of the additional phenylalanine group before the terminal K. This additional F increases the overall hydrophobicity of the peptides and the intermolecular forces (H-bond number) between peptides, resulting in more rigid fibres with higher moduli, without affecting the exponent. For the F hydrogels, significantly higher values of $G^{\prime}$ were observed (due to higher crosslinking levels) with a smaller exponent, as predicted by the models.
[…]
Figure 4: (a)-(c) TEM images of three SAPHs with 10$\times$ dilution in 57000$\times$ magnification, and their chemical structure (inset). (d) Rheology test results for the SAPHs; $G^{\prime}$, the storage modulus, versus peptide concentration (which is a proxy for $\rho l$ in the models due to the one-to-one relationship between the mass and length of a peptide fibre).
In conclusion, to investigate the effects of network structure on material properties, new networks based on Mikado networks were created and studied at low (near percolation) and high densities. Their behaviours were shown to be affected by localised modifications, which represent the existence of local interactions between the fibres when the network is forming. We found that increasing spatial dispersion, and modifying orientation such that fibres have a greater tendency to intersect, decreases the percolation threshold and elastic exponent near the percolation threshold. At high densities, in the non-affine regime, the trends in the model networks’ shear moduli agreed with experiments we conducted on real SAPHs. Finally, we showed that the Poisson’s ratio depends on both the deformation regime and the density. The reason for this lies in how the networks absorb energy at different densities.